OpenStudy (anonymous):

What is the factored form of the expression? d^2-13d+30 a. (d+10)(d+3) b. (d+10)(d-3) c. (d-10)(d+3) d. (d-10)(d-3)

OpenStudy (adamk):

FOIL out all the answer choices if you have to.

OpenStudy (anonymous):

\[d^2-13d+30 =\\d^2-10d-3d+30=\\d(d-10)-3(d-10)=\\(d-3)(d-10)\Rightarrow D\]